# Mixtures of Experts Unlock Parameter Scaling for Deep RL

### v1

@@ Abstract
The recent rapid progress in (self) supervised learning models is in large part predicted by empirical scaling laws: a model’s performance scales proportionally to its size. Analogous scaling laws remain elusive for reinforcement learning domains, however, where increasing the parameter count of a model often hurts its final performance. In this paper, we demonstrate that incorporating Mixture-of-Expert (MoE) modules, and in particular Soft MoEs (Puigcerver et al., 2023), into value-based networks results in more parameter-scalable models, evidenced by substantial performance increases across a variety of training regimes and model sizes. This work thus provides strong empirical evidence towards developing scaling laws for reinforcement learning.
@@ 2.1 Reinforcement Learning
When $\mathcal{X}$ is very large (or infinite), function approximators are used to express $Q$, e.g., DQN by Mnih et al. (2015) uses neural networks with parameters $\theta$, denoted as $Q_{\theta}\approx Q$. The original architecture used in DQN, hereafter referred to as the CNN architecture, comprises of 3 convolutional layers followed by 2 dense layers, with ReLu nonlinearities (Fukushima, 1969) between each pair of layers. In this work, we mostly use the newer Impala architecture (Espeholt et al., 2018), but will provide a comparison with the original CNN architecture. DQN also used a replay buffer: a (finite) memory where an agent stores transitions received while interacting with the environment, and samples mini-batches from to compute gradient updates. The replay ratio is defined as the ratio of gradient updates to environment interactions, and plays a role in our analyses below.
@@ 2.2 Mixtures of Experts
Hard assignments of tokens to experts introduce a number of challenges such as training instabilities, dropping of tokens, and difficulties in scaling the number of experts (Fedus et al., 2022; Puigcerver et al., 2023). To address some of these challenges, Puigcerver et al. (2023) introduced Soft MoE, which is a fully differentiable soft assignment of tokens-to-experts, replacing router-based hard token assignments. Soft assignment is achieved by computing (learned) mixes of per-token weightings for each expert, and averaging their outputs. Following the notation of Puigcerver et al. (2023), let us define the input tokens as $\mathbf{X}\in\mathbb{R}^{m\times d}$, where $m$ is the number of $d$-dimensional tokens. A Soft MoE layer applies a set of $n$ experts on individual tokens, $\left\{f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\right\}_{1:n}$. Each expert has $p$ input- and output-slots, represented respectively by a $d$-dimensional vector of parameters. We denote these parameters by $\boldsymbol{\Phi}\in\mathbb{R}^{d\times(n\cdot p)}$.
@@
Note how $\mathbf{D}$ and $\mathbf{C}$ are learned only through $\boldsymbol{\Phi}$, which we will use in our analysis. Puigcerver et al. (2023) results suggest that Soft MoE achieves a better trade-off between accuracy and computational cost compared to other MoE methods.
@@ 3 Mixture of Experts for Deep RL
Following the predominant usage of MoEs to replace dense feed-forward layers, we replace the penultimate layer in our networks with an MoE module, where each expert has the same dimensionality as the original dense layer. Thus, we are effectively widening the penultimate layer’s dimensionality by a factor equal to the number of experts. Figure 2 illustrates our deep RL MoE architecture. We discuss some alternatives in Section 5.3.
@@ 4 Empirical evaluation
We focus our investigation on DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018), two value-based agents that have formed the basis of a large swath of modern deep RL research. Recent work has demonstrated that using the ResNet architecture  (Espeholt et al., 2018) instead of the original CNN architecture (Mnih et al., 2015) yields strong empirical improvements (Graesser et al., 2022; Schwarzer et al., 2023), so we conduct most of our experiments with this architecture. As in the original papers, we evaluate on 20 games from the Arcade Learning Environment (ALE), a collection of a diverse and challenging pixel-based environments (Bellemare et al., 2013).
Our implementation, included with this submission, is built on the Dopamine library (Castro et al., 2018), which adds stochasticity (via sticky actions) to the ALE (Machado et al., 2018). We use the recommendations from Agarwal et al. (2021) for statistically-robust performance evaluations, in particular focusing on interquartile mean (IQM). Every experiment was run for 200M environment steps, unless reported otherwise, with 5 independent seeds, and we report 95% stratified bootstrap confidence intervals. All experiments were run on NVIDIA Tesla P100 GPUs, and each took on average 4 days to complete.
@@ 5 Future directions
Figure 12: Normalized performance across 26 Atari games for DrQ($\epsilon$) (left) and DER (right), with the ResNet architecture (Espeholt et al., 2018) and 8 experts. Soft MoE not only remains generally stable with more training, but also attains higher final performance. We report interquantile mean performance with error bars indicating 95% confidence intervals.
@@ 5.2 Agent Variants For Low-Data Regimes
DQN and Rainbow were both developed for a training regime where agents can take millions of steps in their environments. Kaiser et al. (2020) introduced the 100k benchmark, which evaluates agents on a much smaller data regime (100k interactions) on 26 games. We evaluate the performance on two popular agents for this regime: DrQ($\epsilon$), an agent based on DQN (Yarats et al., 2021; Agarwal et al., 2021), and DER (Van Hasselt et al., 2019), an agent based on Rainbow. When trained for 100k environment interactions we saw no real difference between the different variants, suggesting that the benefits of MoEs, at least in our current setup, only arise when trained for a significant amount of interactions. However, when trained for 50M steps we do see gains in both agents, in particular with DER (Figure 12). The gains we observe in this setting are consistent with what we have observed so far, given that DrQ($\epsilon$) and DER are based on DQN and Rainbow, respectively.
@@ 5.3 Expert Variants
Figure 13: Normalized performance over 20 games for expert variants with DQN (left) and Rainbow (right), also investigating the use of the normalization of Puigcerver et al. (2023). We report interquantile mean performance with errorbars indicating 95% confidence intervals.
@@
Figure 13 summarizes our results, from which we can draw a number of conclusions. First, these experiments confirm our intuition that not including regularization in the setup used in the majority of this paper performs best. Second, consistent with Puigcerver et al. (2023), normalization seems to help with Big experts. This is particularly noticeable in DQN, where it even surpasses the performance of the Regular setup; the difference between the two is most likely due to the last layer being shared (as in Rainbow) or non-shared (as in DQN). Finally, the using separate MoE modules for each layer (All) seems to not provide many gains, especially when coupled with normalization, where the agents are completely unable to learn.
@@ 6 Related Work
MoEs were first proposed by Jacobs et al. (1991) and have recently helped scaling language models up to trillions of parameters thanks to their modular nature, facilitating distributed training, and improved parameter efficiency at inference (Lepikhin et al., 2020; Fedus et al., 2022). MoEs are also widely studied in computer vision Wang et al. (2020); Yang et al. (2019); Abbas and Andreopoulos (2020); Pavlitskaya et al. (2020), where they enable scaling vision transformers to billions of parameters while reducing the inference computation cost by half (Riquelme et al., 2021), help with vision-based continual learning (Lee et al., 2019), and multi-task problems (Fan et al., 2022). MoEs also help performance in transfer- and multi-task learning settings, e.g. by specializing experts to sub-problems (Puigcerver et al., 2020; Chen et al., 2023; Ye and Xu, 2023) or by addressing statistical performance issues of routers (Hazimeh et al., 2021). There have been few works exploring MoEs in RL for single (Ren et al., 2021; Akrour et al., 2022) and multi-task learning (Hendawy et al., 2024). However their definition and usage of “mixtures-of-experts” is somewhat different than ours, focusing more on orthogonal, probabilistic, and interpretable MoEs.
@@
Lack of parameter scalability in deep RL can be partly explained by a lack of parameter efficiency. Parameter count cannot be scaled efficiently if those parameters are not used effectively. Recent work shows that networks in RL under-utilize their parameters. Sokar et al. (2023) demonstrates that networks suffer from an increasing number of inactive neurons throughout online training. Similar behavior is observed by Gulcehre et al. (2022) in offline RL. Arnob et al. (2021) shows that 95% of the network parameters can be pruned at initialization in offline RL without loss in performance. Numerous works demonstrate that periodic resetting of the network weights improves performance (Nikishin et al., 2022; Dohare et al., 2021; Sokar et al., 2023; D’Oro et al., 2022; Igl et al., 2020; Schwarzer et al., 2023). Another line of research demonstrates that RL networks can be trained with a high sparsity level ($\sim$90%) without loss in performance Tan et al. (2022); Sokar et al. (2022); Graesser et al. (2022). These observations call for techniques to better utilize the network parameters in RL training, such as using MoEs, which we show decreases dormant neurons drastically over multiple tasks and architectures. To enable scaling networks in deep RL, prior works focus on algorithmic methods (Farebrother et al., 2022; Taiga et al., 2022; Schwarzer et al., 2023). In contrast, we focus on alternative network topologies to enable robustness towards scaling.
@@ 7 Discussion and Conclusion
Even in the narrow setting we have focused on (replacing the penultimate layer of value-based agents with MoEs for off-policy learning in single-task settings), there are many open questions that can further increase the benefits of MoEs: different values of $k$ for Top1-MoEs, different tokenization choices, using different learning rates (and perhaps optimizers) for routers, among others. Of course, expanding beyond the ALE could provide more comprehensive results and insights, potentially at a fraction of the computational expense (Ceron and Castro, 2021).
@@ Acknowledgements
The authors would like to thank Gheorghe Comanici, Owen He, Alex Muzio, Adrien Ali Taïga, Rishabh Agarwal, Hugo Larochelle, and the rest of the Google DeepMind Montreal team for valuable discussions during the preparation of this work; Gheorghe deserves a special mention for providing us valuable feedback on an early draft of the paper.
Appendix A Extra results

### v2

@@ Abstract
The recent rapid progress in (self) supervised learning models is in large part predicted by empirical scaling laws: a model’s performance scales proportionally to its size. Analogous scaling laws remain elusive for reinforcement learning domains, however, where increasing the parameter count of a model often hurts its final performance. In this paper, we demonstrate that incorporating Mixture-of-Expert (MoE) modules, and in particular Soft MoEs (Puigcerver et al., 2023), into value-based networks results in more parameter-scalable models, evidenced by substantial performance increases across a variety of training regimes and model sizes. This work thus provides strong empirical evidence towards developing scaling laws for reinforcement learning. We make our code publicly available.
Machine Learning, ICML
@@ 2.1 Reinforcement Learning
When $\mathcal{X}$ is very large (or infinite), function approximators are used to express $Q$, e.g., DQN by Mnih et al. (2015) uses neural networks with parameters $\theta$, denoted as $Q_{\theta}\approx Q$. The original architecture used in DQN, hereafter referred to as the CNN architecture, comprises of 3 convolutional layers followed by 2 dense layers, with ReLu nonlinearities (Fukushima, 1969) between each pair of layers. In this work, we mostly use the newer Impala architecture (Espeholt et al., 2018), but will provide a comparison with the original CNN architecture. DQN also used a replay buffer: a (finite) memory where an agent stores transitions received while interacting with the environment, and samples mini-batches from to compute gradient updates. The replay ratio222 In the hyperparameters established in (Mnih et al., 2015), the policy is updated every 4 environment steps collected, resulting in a replay ratio of 0.25. is defined as the ratio of gradient updates to environment interactions, and plays a role in our analyses below.
@@ 2.2 Mixtures of Experts
Such hard assignments of tokens to experts introduce a number of challenges such as training instabilities, dropping of tokens, and difficulties in scaling the number of experts (Fedus et al., 2022; Puigcerver et al., 2023). To address some of these challenges, Puigcerver et al. (2023) introduced Soft MoE, which is a fully differentiable soft assignment of tokens-to-experts, replacing router-based hard token assignments.
Soft assignment is achieved by computing (learned) mixes of per-token weightings for each expert, and averaging their outputs. Following the notation of Puigcerver et al. (2023), let us define the input tokens as $\mathbf{X}\in\mathbb{R}^{m\times d}$, where $m$ is the number of $d$-dimensional tokens. A Soft MoE layer applies a set of $n$ experts on individual tokens, $\left\{f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\right\}_{1:n}$. Each expert has $p$ input- and output-slots, represented respectively by a $d$-dimensional vector of parameters. We denote these parameters by $\boldsymbol{\Phi}\in\mathbb{R}^{d\times(n\cdot p)}$.
@@
Note how $\mathbf{D}$ and $\mathbf{C}$ are learned only through $\boldsymbol{\Phi}$, which we will use in our analysis. The results of Puigcerver et al. (2023) suggest that Soft MoE achieves a better trade-off between accuracy and computational cost compared to other MoE methods.
@@ 3 Mixture of Experts for Deep RL
Following the predominant usage of MoEs to replace dense feed-forward layers (Shazeer et al., 2017; Fedus et al., 2022; Gale et al., 2023), we replace the penultimate layer in our networks with an MoE module, where each expert has the same dimensionality as the original dense layer. Thus, we are effectively widening the penultimate layer’s dimensionality by a factor equal to the number of experts. Figure 2 illustrates our deep RL MoE architecture. We discuss some alternatives in Section 5.3.
@@ 4 Empirical evaluation
We focus our investigation on DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018), two value-based agents that have formed the basis of a large swath of modern deep RL research. Recent work has demonstrated that using the ResNet architecture  (Espeholt et al., 2018) instead of the original CNN architecture (Mnih et al., 2015) yields strong empirical improvements (Graesser et al., 2022; Schwarzer et al., 2023), so we conduct most of our experiments with this architecture. As in the original papers, we evaluate on 20 games from the Arcade Learning Environment (ALE), a collection of a diverse and challenging pixel-based environments (Bellemare et al., 2013b).
Our implementation, included with this submission, is built on the Dopamine library (Castro et al., 2018) [^3] Dopamine code available at https://github.com/google/dopamine., which adds stochasticity (via sticky actions) to the ALE (Machado et al., 2018). We use the recommendations from Agarwal et al. (2021) for statistically-robust performance evaluations, in particular focusing on interquartile mean (IQM). Every experiment was run for 200M environment steps, unless reported otherwise, with 5 independent seeds, and we report 95% stratified bootstrap confidence intervals. All experiments were run on NVIDIA Tesla P100 GPUs, and each took on average 4 days to complete.
@@ 5.2 Agent Variants For Low-Data Regimes
DQN and Rainbow were both developed for a training regime where agents can take millions of steps in their environments. Kaiser et al. (2020) introduced the 100k [^4] Here, 100k refers to agent steps, or 400k environment frames, due to skipping frames in the standard training setup. benchmark, which evaluates agents on a much smaller data regime (100k interactions) on 26 games. We evaluate the performance on two popular agents for this regime: DrQ($\epsilon$), an agent based on DQN (Yarats et al., 2021; Agarwal et al., 2021), and DER (Van Hasselt et al., 2019), an agent based on Rainbow. When trained for 100k environment interactions we saw no real difference between the different variants, suggesting that the benefits of MoEs, at least in our current setup, only arise when trained for a significant amount of interactions. However, when trained for 50M steps we do see gains in both agents, in particular with DER (Figure 12). The gains we observe in this setting are consistent with what we have observed so far, given that DrQ($\epsilon$) and DER are based on DQN and Rainbow, respectively.
@@ 5.3 Expert Variants
Figure 13 summarizes our results, from which we can draw a number of conclusions. First, these experiments confirm our intuition that not including regularization in the setup used in the majority of this paper performs best. Second, consistent with Puigcerver et al. (2023), normalization seems to help with Big experts. This is particularly noticeable in DQN, where it even surpasses the performance of the Regular setup; the difference between the two is most likely due to the last layer being shared (as in Rainbow) or non-shared (as in DQN). Finally, using separate MoE modules for each layer (All) seems to not provide many gains, especially when coupled with normalization, where the agents are completely unable to learn.
In summary, our results suggest there may be promise in exploring alternative architectures for use with mixtures of experts.
@@ 6 Related Work
MoEs were first proposed by Jacobs et al. (1991) and have recently helped scaling language models up to trillions of parameters thanks to their modular nature, facilitating distributed training, and improved parameter efficiency at inference (Lepikhin et al., 2020; Fedus et al., 2022). MoEs are also widely studied in computer vision (Wang et al., 2020; Yang et al., 2019; Abbas & Andreopoulos, 2020; Pavlitskaya et al., 2020), where they enable scaling vision transformers to billions of parameters while reducing the inference computation cost by half (Riquelme et al., 2021), help with vision-based continual learning (Lee et al., 2019), and multi-task problems (Fan et al., 2022). MoEs also help performance in transfer- and multi-task learning settings, e.g. by specializing experts to sub-problems (Puigcerver et al., 2020; Chen et al., 2023; Ye & Xu, 2023) or by addressing statistical performance issues of routers (Hazimeh et al., 2021). There have been few works exploring MoEs in RL for single (Ren et al., 2021; Akrour et al., 2022) and multi-task learning (Hendawy et al., 2024). However their definition and usage of “mixtures-of-experts” is somewhat different than ours, focusing more on orthogonal, probabilistic, and interpretable MoEs.
@@
Lack of parameter scalability in deep RL can be partly explained by a lack of parameter efficiency. Parameter count cannot be scaled efficiently if those parameters are not used effectively. Recent work shows that networks in RL under-utilize their parameters. Sokar et al. (2023) demonstrates that networks suffer from an increasing number of inactive neurons throughout online training. Similar behavior is observed by Gulcehre et al. (2022) in offline RL. Arnob et al. (2021) shows that 95% of the network parameters can be pruned at initialization in offline RL without loss in performance. Numerous works demonstrate that periodic resetting of the network weights improves performance (Nikishin et al., 2022; Dohare et al., 2021; Sokar et al., 2023; D’Oro et al., 2022; Igl et al., 2020; Schwarzer et al., 2023).
Another line of research demonstrates that RL networks can be trained with a high sparsity level ($\sim$90%) without loss in performance (Tan et al., 2022; Sokar et al., 2022; Graesser et al., 2022; Obando-Ceron et al., 2024). These observations call for techniques to better utilize the network parameters in RL training, such as using MoEs, which we show decreases dormant neurons drastically over multiple tasks and architectures. To enable scaling networks in deep RL, prior works focus on algorithmic methods (Farebrother et al., 2022; Taiga et al., 2022; Schwarzer et al., 2023; Farebrother et al., 2024). In contrast, we focus on alternative network topologies to enable robustness towards scaling.
@@ 7 Discussion and Conclusion
Figure 12: Normalized performance across 26 Atari games for DrQ($\epsilon$) (left) and DER (right), with the ResNet architecture (Espeholt et al., 2018) and 8 experts (see Figure 16, for 4 experts). Soft MoE not only remains generally stable with more training, but also attains higher final performance. We report interquantile mean performance with error bars indicating 95% confidence intervals.
@@
Even in the narrow setting we have focused on (replacing the penultimate layer of value-based agents with MoEs for off-policy learning in single-task settings), there are many open questions that can further increase the benefits of MoEs: different values of $k$ for Top1-MoEs, different tokenization choices, using different learning rates (and perhaps optimizers) for routers, among others. Of course, expanding beyond the ALE could provide more comprehensive results and insights, potentially at a fraction of the computational expense (Ceron & Castro, 2021).
Figure 13: Normalized performance over 20 games for expert variants with DQN (left) and Rainbow (right), also investigating the use of the normalization of Puigcerver et al. (2023). We report interquantile mean performance with shaded areas indicating 95% confidence intervals.
@@ Acknowledgements
The authors would like to thank Gheorghe Comanici, Owen He, Alex Muzio, Adrien Ali Taïga, Rishabh Agarwal, Hugo Larochelle, Ayoub Echchahed, and the rest of the Google DeepMind Montreal team for valuable discussions during the preparation of this work; Gheorghe deserves a special mention for providing us valuable feedback on an early draft of the paper. We thank the anonymous reviewers for their valuable help in improving our manuscript. We would also like to thank the Python community (Van Rossum & Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007) and JAX (Bradbury et al., 2018).
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning, and reinforcement learning in particular. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Experimental details
Unless otherwise specified, in all experiments below we report the interquantile mean after 40 million environment steps; error bars indicate 95% stratified bootstrap confidence intervals (Agarwal et al., 2021). Most of our experiments were run with 20 games from the ALE suite (Bellemare et al., 2013a), as suggested by Fedus et al. (2020). However, for the Atari 100k agents (subsection 5.2), we used the standard set of 26 games (Kaiser et al., 2020) to be consistent with the benchmark. Finally, we also ran some experiments with the full set of 60 games. The specific games are detailed below.
20 game subset: AirRaid, Asterix, Asteroids, Bowling, Breakout, DemonAttack, Freeway, Gravitar, Jamesbond, MontezumaRevenge, MsPacman, Pong, PrivateEye, Qbert, Seaquest, SpaceInvaders, Venture, WizardOfWor, YarsRevenge, Zaxxon.
26 game subset: Alien, Amidar, Assault, Asterix, BankHeist, BattleZone, Boxing, Breakout, ChopperCommand, CrazyClimber, DemonAttack, Freeway, Frostbite, Gopher, Hero, Jamesbond, Kangaroo, Krull, KungFuMaster, MsPacman, Pong, PrivateEye, Qbert, RoadRunner, Seaquest, UpNDown.
60 game set: The 26 games above in addition to: AirRaid, Asteroids, Atlantis, BeamRider, Berzerk, Bowling, Carnival, Centipede, DoubleDunk, ElevatorAction, Enduro, FishingDerby, Gravitar, IceHockey, JourneyEscape, MontezumaRevenge, NameThisGame, Phoenix, Pitfall, Pooyan, Riverraid, Robotank, Skiing, Solaris, SpaceInvaders, StarGunner, Tennis, TimePilot, Tutankham, Venture, VideoPinball, WizardOfWor, YarsRevenge, Zaxxon.
Appendix B Hyper-parameters list
Default hyper-parameter settings for DER (Van Hasselt et al., 2019) and DrQ($\epsilon$) (Kaiser et al., 2020; Agarwal et al., 2021). Table 1 shows the default values for each hyper-parameter across all the Atari games.
Table 1: Default hyper-parameters setting for DER and DrQ($\epsilon$) agents.
Atari
Hyper-parameter DER DrQ($\epsilon$)
Adam’s($\epsilon$) 0.00015 0.00015
Batch Size 32 32
Conv. Activation Function ReLU ReLU
Convolutional Width 1 1
Dense Activation Function ReLU ReLU
Dense Width 512 512
Normalization None None
Discount Factor 0.99 0.99
Exploration $\epsilon$ 0.00015 0.00015
Learning Rate 0.01 0.01
Minimum Replay History 1600 1600
Number of Atoms 51 0
Number of Convolutional Layers 3 3
Number of Dense Layers 2 2
Replay Capacity 1000000 1000000
Reward Clipping True True
Update Horizon 10 10
Update Period 1 1
Weight Decay 0 0
Sticky Actions False False
Default hyper-parameter settings for DQN (Mnih et al., 2015), Rainbow (Hessel et al., 2018). Table 2 shows the default values for each hyper-parameter across all the Atari games.
Table 2: Default hyper-parameters setting for DQN, Rainbow, IQN, Munchausen-IQN agents.
Atari
Hyper-parameter DQN Rainbow IQN M-IQN
Adam’s($\epsilon$) 1.5e-4 1.5e-4 1.5e-4 1.5e-4
Batch Size 32 32 32 32
Conv. Activation Function ReLU ReLU ReLU ReLU
Convolutional Width 1 1 1 1
Dense Activation Function ReLU ReLU ReLU ReLU
Dense Width 512 512 512 512
Normalization None None None None
Discount Factor 0.99 0.99 0.99 0.99
Exploration $\epsilon$ 0.01 0.01 0.01 0.01
Learning Rate 6.25e-5 6.25e-5 0.00005 0.00005
Minimum Replay History 20000 20000 20000 20000
Number of Atoms 0 51 - -
Kappa - - 1.0 1.0
Num tau samples - - 64 64
Num tau prime samples - - 64 64
Num quantile samples - - 32 32
Number of Convolutional Layers 3 3 3 3
Number of Dense Layers 2 2 2 2
Replay Capacity 1000000 1000000 1000000 1000000
Reward Clipping True True True True
Update Horizon 1 3 3 3
Update Period 4 4 4 4
Weight Decay 0 0 0 0
Sticky Actions True True True True
Tau - - 0 0.03
Default hyper-parameter settings for CQL (Kumar et al., 2020) and CQL+C51 (Kumar et al., 2022) offline agents. Table 3 shows the default values for each hyper-parameter across all the Atari games.
Table 3: Default hyper-parameters setting for CQL and CQL+C51 agents.
Atari
Hyper-parameter CQL CQL+C51
Adam’s($\epsilon$) 0.0003125 0.00015
Batch Size 32 32
Conv. Activation Function ReLU ReLU
Convolutional Width 1 1
Dense Activation Function ReLU ReLU
Normalization None None
Dense Width 512 512
Discount Factor 0.99 0.99
Learning Rate 0.00005 0.0000625
Number of Atoms 0 51
Number of Convolutional Layers 3 3
Number of Dense Layers 2 2
Fixed Replay Capacity 2,500,000 steps 2,500,000 steps
Reward Clipping True True
Update Horizon 1 3
Update Period 1 1
Weight Decay 0 0
Replay Scheme Uniform Uniform
Dueling True True
Double DQN True True
CQL coef 0.1 0.1
Default hyper-parameter settings for CNN architecture (Mnih et al., 2015) and Impala-based ResNet (Espeholt et al., 2018) Table 4 shows the default values for each hyper-parameter across all the Atari games.
Table 4: Default hyper-parameters for neural networks.
Atari
Hyper-parameter CNN architecture (Mnih et al., 2015) Impala-based ResNet (Espeholt et al., 2018)
Observation down-sampling (84, 84) (84, 84)
Frames stacked 4 4
Q-network (channels) 32, 64, 64 32, 64, 64
Q-network (filter size) 8 x 8, 4 x 4, 3 x 3 8 x 8, 4 x 4, 3 x 3
Q-network (stride) 4, 2, 1 4, 2, 1
Num blocks - 2
Use max pooling False True
Skip connections False True
Hardware Tesla P100 GPU Tesla P100 GPU
Appendix C Extra results